Calculate the amount of Retin-A® 0.05% Cream in grams required to prepare 3 pounds of RHHA cream that has the following formula.
Retin-A® 0.05% Cream 30 parts
Hydrocortisone 2.5% Cream 25 parts
Hydroquinone Powder 32 parts
Vitamin C 6 parts


      
A. 87.9 g
       
B. 806.5 g


       
C. 468.6 g
       
D. 366.1 g
       
E. 439.4 g
Answer #1

Mass of RHHA cream in grams = mass in lb * 453.592 = 3 * 453.592 = 1360.78 g

Mass of Retin - A = parts of Retin A / total parts * mass in g = 30 / (30 + 25 + 32 + 6) * 1360.78 = 439.4 g